OpenStudy (anonymous):

The equation of a line is y=-1/2x-1. what is the equation of the line that is perpendicular to the first line and passes through the point (2,-5)?

OpenStudy (anonymous):

You will have follow this steps: To find the slope of the perpendicular: Slope=opposite inverse of the perpendicular's slope. Then you will use this: y−y1=m(x−x1) Where m is the slope and (x1, y1) is the point Please, tell us what are you getting.

OpenStudy (anonymous):

I still dont get it. Can you help me some more? please

OpenStudy (anonymous):

We have that the slope of the original line is: -1/2 Which is the opposite inverse of -1/2?

OpenStudy (anonymous):

@NaeNae16 Are you still with us? :)

OpenStudy (anonymous):

but im still not sure what to do

OpenStudy (anonymous):

The opposite inverse of -1/2=2 So the slope of the line that we are looking for is m=2 Now you have to use this equation: y-y1=m*(x-x1) and using this: (x1,y1)=(2,-5) Come on, give it a try!

OpenStudy (anonymous):

yes I get it Thanks :)